# LRFusionPR: A Polar BEV-Based LiDAR-Radar Fusion Network for Place Recognition

### v2

@@ Abstract
In autonomous driving, place recognition is critical for global localization in GPS-denied environments. LiDAR and radar-based place recognition methods have garnered increasing attention, as LiDAR provides precise ranging, whereas radar excels in adverse weather resilience. However, effectively leveraging LiDAR-radar fusion for place recognition remains challenging. The noisy and sparse nature of radar data limits its potential to further improve recognition accuracy. In addition, heterogeneous radar configurations complicate the development of unified cross-modality fusion frameworks. In this paper, we propose LRFusionPR, which improves recognition accuracy and robustness by fusing LiDAR with either single-chip or scanning radar. Technically, a dual-branch network is proposed to fuse different modalities within the unified polar coordinate bird’s eye view (BEV) representation. In the fusion branch, cross-attention is utilized to perform cross-modality feature interactions. The knowledge from the fusion branch is simultaneously transferred to the distillation branch, which takes radar as its only input to further improve the robustness. Ultimately, the descriptors from both branches are concatenated, producing the multimodal global descriptor for place retrieval. Extensive evaluations on multiple datasets demonstrate that our LRFusionPR achieves accurate place recognition, while maintaining robustness under varying weather conditions. Our open-source code will be released at https://github.com/QiZS-BIT/LRFusionPR.
@@ I Introduction
Place recognition is a localization technique that uses perceptual data to identify whether a vehicle is revisiting a location. It works by encoding scenes into compact global descriptors, which are then matched against a database using similarity metrics [1, 2, 3, 4]. This capability enables global localization in GPS-denied environments and supports loop closure detection in SLAM systems. In autonomous driving applications, LiDAR place recognition (LPR) has attracted increasing attention [2, 5, 6, 7]. LiDAR captures precise structural information and exhibits robustness under varying illumination conditions. These attributes endow LPR with significant potential.
Recent studies have focused on integrating LiDAR with other sensors, leveraging their complementary strengths to enhance recognition accuracy. The most prevalent multimodal place recognition (MPR) methods combine LiDAR with cameras [8, 9, 10, 11], which provide rich semantic and textural information. However, both LiDAR and camera share a common drawback, exhibiting suboptimal performance under adverse weather conditions. For instance, under the influence of fog, the maximum perception range of LiDAR point clouds decreases and noise emerges, while the appearance of images undergoes significant alterations.
@@
To address these limitations, radar stands out as a promising alternative sensor. Although radar offers lower data quality, its exceptional signal penetration capability provides superior weather robustness. However, how to effectively perform LiDAR-radar fusion for place recognition remains an open problem. The main challenges arise from two aspects. Firstly, the two most common types of radar in radar place recognition (RPR) differ in their perceptual formats. Single-chip radar generates sparse point clouds with velocity, while scanning radar outputs polar intensity maps. This domain gap limits existing RPR methods to specific radar types [3, 12, 4, 13]. Secondly, radar data exhibits sparse and noisy characteristics, which complicates the extraction of discriminative representations.
Despite these technical hurdles, recent advances have paved the way for effective LiDAR-radar fusion in place recognition. BEVPlace [14] validates the efficacy of BEV-formatted LiDAR data for place recognition, thus demonstrating the feasibility of unifying heterogeneous sensor formats. Bi-LRFusion [15] and RadarDistill [16] show the effectiveness of exploiting the feature correspondences between LiDAR and radar in 3D object detection tasks.
In this paper, we propose a LiDAR-radar fusion-based MPR method dubbed LRFusionPR, as shown in Fig. 1. Firstly, LiDAR and radar data format are unified through polar orthogonal BEV projection. Subsequently, a dual-branch network exploits cross-modality feature correlations within polar BEV representations. The fusion branch enhances recognition accuracy through cross-attention, integrating spatially aligned LiDAR and radar features. The distillation branch operates on radar-only inputs, fully unleashing the radar-specific robustness. Furthermore, to enhance robustness while retaining accuracy, we distill the knowledge from fusion branch into distillation branch. Ultimately, descriptors from both branches are concatenated into a global descriptor for robust and accurate place recognition.
@@
  • We propose a polar BEV-based LiDAR-radar fusion network, LRFusionPR, to achieve accurate and robust place recognition in autonomous driving scenarios.
  • We propose the LiDAR-radar feature fusion mechanism based on a dual-branch network. Cross-attention and distillation are synergistically integrated to preserve the specific advantages of both LiDAR and radar while performing feature fusion.
  • Extensive evaluations on three datasets and adverse weather conditions demonstrate our method’s performance in both accuracy and robustness.
@@ II Related Work
Place recognition has been extensively studied in computer vision and robotics, with numerous works proposed on visual place recognition (VPR). We recommend referring to [17] for an in-depth overview of VPR. Here we focus specifically on discussing LiDAR place recognition (LPR) and radar place recognition (RPR).
@@ II-A LiDAR Place Recognition
LiDAR’s robustness to illumination and seasonal variations makes LPR suitable for autonomous driving scenarios. Scan Context [18] construct descriptors by encoding the distribution of LiDAR point clouds. PointNetVLAD [2] takes downsampled point cloud submaps as input to perform point-wise feature extraction and aggregation. Building on this, LPDNet [19] further combines hand-crafted local spatial features with deep learning. OverlapNet [20], OverlapTransformer [6] and SeqOT [21] project point clouds into range images, enabling efficient feature extraction. EgoNN [5] utilizes 3D sparse convolution to extract spatial features from sparsely distributed LiDAR point clouds. BEVPlace [14] and BEVPlace++ [7] project point clouds into BEV images to extract discriminative descriptors.
More recently, the integration of LiDAR and camera data has emerged as a paradigm for multimodal place recognition (MPR). MinkLoc++ [8] and AdaFusion [22] enhance recognition accuracy by fusing descriptors from different modalities. LCPR [9] and EINet [10] exploit cross-modality correspondence through interactive feature fusion.
@@ II-B Radar Place Recognition
While the radar shows a lower data quality than LiDAR, its unique signal penetration capability provides critical operational advantages in adverse weather conditions. These advantages have made RPR a viable solution. Single-chip radar and scanning radar are the two main radar types that have been extensively studied in RPR. Due to its comparatively dense and accurate nature, scanning radars has attracted considerable attention. KidnappedRadar [3] achieves robust place recognition by extracting rotation-invariant features. RaPlace [23] proposes a non-learning method utilizing the Radon transform. Compared to scanning radar, single-chip radar data is sparser and noisier, making it difficult to generate discriminative representations. AutoPlace [4] improves the recognition accuracy by exploiting sequential information. SPR [13] achieves single frame-based place recognition by combining an importance estimation module with RCS data. CRPlace [24] proposes an MPR approach that integrates single-chip radar with camera.
@@ II-C LiDAR-Radar Fusion
Recently, the combination of LiDAR and radar has attracted interest in 3D object detection. LiDAR and cameras are prone to degradation in adverse weather, resulting in poor robustness. In contrast, radar’s weather resilience and LiDAR’s high resolution form a strong complement, demonstrating the potential for both robustness and precision. MVDNet [25] integrates LiDAR and radar into a unified framework to improve detection robustness. ST-MVDNet [26] proposes a distillation framework to address the problem of sensor failures. Bi-LRFusion [15] introduces a two-stage fusion mechanism to address the inherent weakness of radar data. RadarDistill [16] employs a knowledge distillation approach that transfers LiDAR-derived features to improve the performance of the radar-only detector.
However, how to boost place recognition performance through LiDAR-radar fusion has not been fully explored. To address this, we propose a novel polar BEV-based MPR method, dubbed LRFusionPR. A dual-branch network is proposed to fully leverage the advantages of both LiDAR and radar, enabling accurate and robust place recognition.
@@ III-A LRFusionPR Architecture
Fig. 2 illustrates the architecture of our proposed LRFusionPR, which fuses LiDAR with single-chip or scanning radar for accurate and robust place retrieval. We first unify multimodal data into the polar BEV representation. This geometric transformation standardizes LiDAR and heterogeneous radar data, effectively eliminating their discrepancies. Using two ResNet backbones [27] of different depths, we then extract spatially aligned features from polar BEVs. Subsequently, the multimodal features are fed into a dual-branch network, consisting of the fusion branch and the distillation branch. The main role of the fusion branch is to fully exploit the cross-modality feature correlations. Due to the inherent sparsity and noise of radar data, direct integration of radar features into a LiDAR-centric network is not optimal [15]. Therefore, we propose a polar cross-attention module to perform deep feature interaction between modalities. The discriminability of radar features is improved under the guidance of LiDAR cues. However, performing feature fusion alone cannot fully exploit the radar-specific robustness. To address this, we introduce the distillation branch. This component processes radar data exclusively to generate descriptors, fully exploiting the inherent strengths of radar. We further propose using the structure-aware loss [28] to guide knowledge transfer from the fusion branch to the distillation branch. This mechanism increases the discriminability of the radar-based descriptors, enhancing the robustness and recognition accuracy. Ultimately, the descriptors from both branches are concatenated to achieve accurate and robust place recognition.
@@ III-B Polar BEV-Based Paradigm
Polar orthogonal projection. LiDAR, single-chip radar, and scanning radar each provide complementary scene understanding through distinct modalities. However, their heterogeneous representations create fundamental domain discrepancies. Thus, unifying the data from different modalities into a common format makes practical sense to ensure effective multimodal feature fusion. Inspired by BEVPlace [14], we unify LiDAR and radar data into polar BEV format via polar orthogonal projection.
Let LiDAR point cloud $\mathcal{P}$ and radar point cloud $\mathcal{R}$ constitute a spatially and temporally aligned multimodal input pair. By performing polar orthogonal projection on $\mathcal{P}$ and $\mathcal{R}$, the LiDAR polar BEV $\mathcal{B}_{\text{P}}$ and the radar polar BEV $\mathcal{B}_{\text{R}}$ are obtained respectively. Using the LiDAR modality as an example, each LiDAR point in the Cartesian coordinate $p_{i}=\left(x_{\text{P}},y_{\text{P}},z_{\text{P}}\right)^{\text{T}}$ is converted to the pixel coordinate $\left(u_{\text{P}},v_{\text{P}}\right)$ of $\mathcal{B}_{\text{P}}$ by:
@@
where $\left(h_{\text{P}},w_{\text{P}}\right)$ represents the height and width of LiDAR polar BEV, and $m$ denotes the maximum perception range. The value of the polar BEV at position $\left(u_{\text{P}},v_{\text{P}}\right)$ is determined by the number of points that fall within the corresponding grid. Similarly, the radar point cloud $\mathcal{R}$ can be transformed into radar polar BEV $\mathcal{B}_{\text{R}}$ with a shape of $\left(h_{\text{R}},w_{\text{R}}\right)$.
Scanning radar data in most datasets [12, 29] are provided as polar intensity maps. We crop polar intensity maps to maximum range $m$, and resize to construct scanning radar BEVs. By unifying different modalities into polar BEV representations, we provide a foundation for cross-modality feature interaction and fusion.
LiDAR/radar encoders. To ensure the lightweight nature of our LRFusionPR, we choose the ResNet18 [27] as the backbone network. To accommodate the resolution discrepancy between LiDAR and radar polar BEVs, we implement asymmetric ResNet encoders. The full ResNet18 is employed as the LiDAR encoder and a shallow invariant (input conv and two ResNet blocks) as the radar encoder. By feeding LiDAR polar BEV $\mathcal{B}_{\text{P}}$ and radar polar BEV $\mathcal{B}_{\text{R}}$ into LiDAR/radar encoders respectively, we can obtain spatially aligned and shape-matched feature maps. We aim to leverage the polar BEV representations to simplify rotation-invariant feature extraction. This approach exploits the geometric property where viewpoint rotations become circular column shifts in polar coordinates. Although spatial downsampling disrupts shift equivariance of CNNs, this limitation can be addressed by the metric learning process. Through training with viewpoint-diverse triplets, the network can regain robustness to input translation. This enables recognition of challenging scenarios like reverse loop closures.
@@ III-C Multimodal Feature Fusion
Polar cross-attention module. The sparsity and noise of radar features often cause simple fusion methods to yield suboptimal results. Inspired by [15], we develop a polar cross-attention module. This module captures cross-modality correlations while enhancing the discriminability of radar fatures through attentive fusion, as detailed in Fig. 3.
Figure 3: The pipeline of the polar cross-attention module. It consists of two cross-attention branches to facilitate extensive feature interaction on the polar feature map, thus exploiting cross-modality correlations.
To facilitate cross-attention, we first align the dimensions of the multimodal polar feature maps. The LiDAR feature is compressed via a convolution layer, reducing its channels from $C_{\text{P}}$ to $C_{\text{R}}$, where $C_{\text{P}}$ is the LiDAR feature channel, and $C_{\text{R}}$ is the radar feature channel. This ensures the LiDAR feature $\mathcal{H}_{\text{P}}$ matches the radar feature $\mathcal{H}_{\text{R}}\in\mathbb{R}^{B\times C_{\text{R}}\times H\times W}$ in dimensions. Next, we flatten both features into sequences $\mathcal{H}_{\text{P}}^{{}^{\prime}}\in\mathbb{R}^{B\times HW\times C_{\text{R}}}$ and $\mathcal{H}_{\text{R}}^{{}^{\prime}}\in\mathbb{R}^{B\times HW\times C_{\text{R}}}$, which are then processed by cross-attention for multimodal fusion.
@@
$\begin{split}\bar{\mathcal{A}}&=\mathrm{LN}\left(\mathrm{Attention}\left(\mathcal{Q}_{\text{R}},\mathcal{K}_{\text{P}},\mathcal{V}_{\text{P}}\right)\right)\\ &+\mathrm{LN}\left(\mathrm{Attention}\left(\mathcal{Q}_{\text{P}},\mathcal{K}_{\text{R}},\mathcal{V}_{\text{R}}\right)\right)\\ \end{split}$ (7)
where $\bar{\mathcal{A}}$ denotes the feature split of $\mathcal{A}$, $\mathrm{LN}(.)$ refers to layer normalization, $\mathcal{Q}_{\text{R}}$, $\mathcal{K}_{\text{R}}$, $\mathcal{V}_{\text{R}}$ are substantially the query, key, and value splits obtained by linearly transforming the splits of $\mathcal{H}_{\text{R}}^{{}^{\prime}}$, and $\mathcal{Q}_{\text{P}}$, $\mathcal{K}_{\text{P}}$, $\mathcal{V}_{\text{P}}$ are those derived by the splits of $\mathcal{H}_{\text{P}}^{{}^{\prime}}$.
Ultimately, the cross-attention enhanced radar feature is concatenated with LiDAR feature to generate the descriptor of the fusion branch. The proposed cross-attention mechanism generates more discriminative feature representations by capturing cross-modal correlations.
Structure-aware distillation. Cross-modality feature fusion facilitates the extraction of discriminative place descriptions. However, in adverse weather conditions, performing feature fusion alone cannot fully exploit the radar-specific robustness. To address this, we propose the radar-exclusive distillation branch to preserve the robustness of radar data. Furthermore, we improve the discriminability of radar features via cross-modality distillation inspired by [26, 16].
The motivation for distillation lies in the fact that the fusion branch exhibits high recognition accuracy but poor robustness, while the the distillation branch exhibits the opposite. We aim to improve the performance of the distillation branch through self-distillation. Knowledge is transferred from the fusion branch to the distillation branch by minimizing the structure-aware loss [28]. The structure-aware loss (formally defined in Sec. III-D) enforces angular constraints on descriptor triplets. This geometric constraint ensures the descriptor triplets maintains its distributions while assimilating structural knowledge from the fusion branch. This enables the distillation branch to learn the discriminability of multimodal fusion features while exploiting the radar-specific robustness. By concatenating the descriptors from both the fusion and distillation branches, place recognition with both accuracy and robustness can be achieved.

### v3

@@ Abstract
In autonomous driving, place recognition is critical for global localization in GPS-denied environments. LiDAR and radar-based place recognition methods have garnered increasing attention, as LiDAR provides precise ranging, whereas radar excels in adverse weather resilience. However, effectively leveraging LiDAR-radar fusion for place recognition remains challenging. The noisy and sparse nature of radar data limits its potential to further improve recognition accuracy. In addition, heterogeneous radar configurations complicate the development of unified cross-modality fusion frameworks. In this paper, we propose LRFusionPR, which improves recognition accuracy and robustness by fusing LiDAR with either single-chip or scanning radar. Technically, a dual-branch network is proposed to fuse different modalities within the unified polar coordinate bird’s eye view (BEV) representation. In the fusion branch, cross-attention is utilized to perform cross-modality feature interactions. The knowledge from the fusion branch is simultaneously transferred to the distillation branch, which takes radar as its only input to further improve the robustness. Ultimately, the descriptors from both branches are concatenated, producing the multimodal global descriptor for place retrieval. Extensive evaluations on multiple datasets demonstrate that our LRFusionPR achieves accurate and rotation-invariant place recognition, while maintaining robustness under varying weather conditions. Our open-source code is available at https://github.com/QiZS-BIT/LRFusionPR.
@@ I Introduction
Place recognition is a localization technique that uses perceptual data to identify whether a vehicle is revisiting a location [1, 2, 3, 4]. This capability enables global localization in GPS-denied environments and supports loop closure detection in SLAM systems. In autonomous driving applications, LiDAR place recognition (LPR) has attracted increasing attention [2, 5, 6, 7]. LiDAR captures precise structural information and exhibits robustness under varying illumination conditions. These attributes endow LPR with significant potential.
Recent studies have focused on integrating LiDAR with other sensors, leveraging their complementary strengths to enhance recognition accuracy. The most prevalent multimodal place recognition (MPR) methods combine LiDAR with cameras [8, 9], which provide rich semantic and textural information. However, LiDAR and camera share a common drawback, exhibiting suboptimal performance under adverse weather conditions.
@@
To address these limitations, radar stands out as a promising alternative sensor. Although radar offers relatively lower data quality, its exceptional signal penetration capability provides superior weather robustness. However, how to effectively perform LiDAR-radar fusion for place recognition remains an open problem. The main challenges arise from two aspects. Firstly, the two most common types of radar in radar place recognition (RPR) differ in their perceptual formats. Single-chip radar generates sparse point clouds with velocity, while scanning radar outputs polar intensity maps. This domain gap limits existing RPR methods to specific radar types [3, 10, 4]. Secondly, radar data exhibits sparse and noisy characteristics, which complicates the extraction of discriminative representations.
In this paper, we propose LRFusionPR, a LiDAR-radar fusion-based MPR method designed to address these challenges, as shown in Fig. 1. The proposed framework demonstrates compatibility with heterogeneous radars, leveraging the individual strengths of both LiDAR and radar to deliver state-of-the-art recognition performance under adverse conditions. To address the challenge arising from heterogeneous data formats, we unify LiDAR and radar data formats through polar orthogonal BEV projection. Moreover, polar BEV facilitates the straightforward and effective extraction of rotation-invariant descriptors, a contrast to Cartesian BEV. Subsequently, a dual-branch network exploits cross-modality feature correlations within polar BEV representations. The fusion branch enhances recognition accuracy through cross-attention, integrating spatially aligned LiDAR and radar features. The distillation branch operates on radar-only inputs, fully unleashing the radar-specific robustness. Furthermore, to facilitate the extraction of discriminative radar features, we distill the knowledge from fusion branch into distillation branch. Ultimately, descriptors from both branches are concatenated for robust and accurate place recognition.
@@
  • We propose a polar BEV-based LiDAR-radar fusion network, LRFusionPR, to achieve accurate and robust place recognition in a rotation-invariant way.
  • We propose the LiDAR-radar feature fusion mechanism based on a dual-branch network. Cross-attention and distillation are synergistically integrated to preserve the specific advantages of both LiDAR and radar.
  • Extensive evaluations on four datasets and adverse weather conditions demonstrate our method’s performance in both accuracy and robustness.
@@ II-A LiDAR Place Recognition
LiDAR’s robustness to illumination and seasonal variations makes LPR suitable for autonomous driving scenarios. Scan Context [11] constructs descriptors by encoding the distribution of LiDAR point clouds. PointNetVLAD [2] takes downsampled point cloud submaps as input to perform point-wise feature extraction. OverlapTransformer [6] project point clouds into range images, enabling efficient feature extraction. MinkLoc3Dv2 [12] and EgoNN [5] utilizes 3D sparse convolution to extract spatial features from sparsely distributed LiDAR point clouds. BEVPlace++ [7] project point clouds into BEV images to extract discriminative descriptors. More recently, the integration of LiDAR and camera has emerged as a paradigm for multimodal place recognition (MPR). MinkLoc++ [8] enhance recognition accuracy by fusing descriptors from different modalities. LCPR [9] exploit cross-modality correspondence through interactive feature fusion.
@@ II-B Radar Place Recognition
While the radar offers relatively lower data quality than LiDAR, its robustness in adverse weather conditions has made RPR a viable solution. Single-chip radar and scanning radar are the two main radar types that have been extensively studied in RPR. Due to its comparatively dense and accurate nature, scanning radars have attracted considerable attention. KidnappedRadar [3] achieves robust place recognition by extracting rotation-invariant features. Off the Radar [13] introduces an uncertainty estimation-based approach for place recognition. Radar SLAM [14] establishes a robust SLAM system capable of operating under all weather conditions. RaPlace [15] proposes a non-learning method utilizing the Radon transform. ReFeree [10] extracts lightweight descriptors using a feature and free space. OpenRadVLAD [16] extracts discriminative descriptors from polar radar scans. Compared to scanning radar, single-chip radar data is sparser and noisier. AutoPlace [4] improves the recognition accuracy by exploiting sequential information. CRPlace [17] proposes an MPR approach that integrates single-chip radar with camera. More recently, 4D RadarPR [18] proposed an RPR framework using 4D radar data.
@@ II-C LiDAR-Radar Fusion
Recently, the combination of LiDAR and radar has attracted interest. LiDAR and cameras are prone to degradation in adverse weather, resulting in poor robustness. In contrast, radar’s weather resilience and LiDAR’s high resolution form a strong complement. Extensive studies in 3D object detection have validated the advantages of LiDAR and radar fusion [19]. In place recognition, some work, such as Radar-to-Lidar [20], has attempted to leverage the complementary strengths of LiDAR and radar through cross-modal alignment.
However, how to boost place recognition performance through LiDAR-radar fusion has not been fully explored. To address this, we propose a novel polar BEV-based MPR method, dubbed LRFusionPR. We aim to propose a framework capable of accommodating both single-chip and scanning radar, enabling accurate and robust place recognition under various weather conditions.
@@ III-A LRFusionPR Architecture
Fig. 2 illustrates the architecture of our proposed LRFusionPR, which fuses LiDAR with single-chip or scanning radar for accurate and robust place retrieval. We first unify multimodal data into the polar BEV representation. Not only does it standardize the data formats between LiDAR and heterogeneous radars, it also provides a data foundation for rotation-invariant place recognition. Using two ResNet backbones [21] of different depths, we then extract spatially aligned features from polar BEVs. Subsequently, the multimodal features are fed into a dual-branch network, consisting of the fusion branch and the distillation branch. The main role of the fusion branch is to fully exploit the cross-modality feature correlations. Due to the inherent sparsity and noise of radar data, direct integration of radar features into a LiDAR-centric network may not be optimal. Therefore, we propose a polar cross-attention module to perform deep feature interaction between modalities. The discriminability of radar features is improved under the guidance of LiDAR cues. However, performing feature fusion alone cannot fully exploit the radar-specific robustness. To address this, we introduce the distillation branch, which processes radar data exclusively to generate robust descriptors. We further propose using the structure-aware loss [22] to guide knowledge transfer from the fusion branch to the distillation branch. This mechanism distills the multimodal fusion knowledge to the distillation branch, thereby increasing the discriminability of radar-based descriptors. Ultimately, the descriptors from both branches are concatenated to achieve accurate and robust place recognition in a rotation-invariant way.
@@ III-B Polar BEV-Based Paradigm
Polar orthogonal projection. We unify LiDAR, single-chip radar, and scanning radar into a unified polar BEV format to address their inherent domain discrepancies.
Let LiDAR point cloud $\mathcal{P}$ and radar point cloud $\mathcal{R}$ constitute a spatially and temporally aligned input pair. By performing polar orthogonal projection on $\mathcal{P}$ and $\mathcal{R}$, the LiDAR polar BEV $\mathcal{B}_{\text{P}}$ and the radar polar BEV $\mathcal{B}_{\text{R}}$ are obtained respectively. Using the LiDAR modality as an example, each LiDAR point in the Cartesian coordinate $p_{i}=\left(x_{\text{P}},y_{\text{P}},z_{\text{P}}\right)^{\text{T}}$ is converted to the pixel coordinate $\left(u_{\text{P}},v_{\text{P}}\right)$ of $\mathcal{B}_{\text{P}}$ by:
@@
where $\left(h_{\text{P}},w_{\text{P}}\right)$ represents the size of LiDAR polar BEV, and $m$ denotes the maximum perception range. To ensure robustness, we represent each cell $\left(u_{\text{P}},v_{\text{P}}\right)$ by its point density, which is calculated from the number of points within it. Similarly, the radar point cloud $\mathcal{R}$ can be transformed into radar polar BEV $\mathcal{B}_{\text{R}}$ with a shape of $\left(h_{\text{R}},w_{\text{R}}\right)$.
Similarly, BEVPlace++ [7] uses point density to construct Cartesian BEV representations. However, as viewpoint rotation causes nonlinear transformations in Cartesian BEV, BEVPlace++ necessitates the use of computationally expensive group convolutions to extract rotation-equivariant features. Thus, we instead select polar BEV. Its inherent property of translating viewpoint rotations into column shifts provides the data foundation for effective and efficient rotation-invariant place recognition (see Sec. IV-F).
LiDAR/radar encoders. To ensure the lightweight nature, we choose the ResNet18 [21] as the backbone. To accommodate the resolution discrepancy between LiDAR and radar polar BEVs, we implement asymmetric ResNet encoders featuring a differing number of ResBlocks. By feeding the polar BEVs into the encoders, we can obtain spatially aligned and shape-matched feature maps.
Our resulting feature maps also exhibit rotation-equivariant properties. Although spatial downsampling breaks our encoders’ strict shift equivariance, making a mathematically explicit demonstration of rotation-equivariance impossible, the nature of polar BEV allows us to overcome this easily through the metric learning process. By training with viewpoint-diverse triplets, the network regains robustness to input translation, thereby achieving rotation-equivariant feature extraction.
Descriptor extraction. We employ NetVLAD [1] to aggregate descriptors from both branches. Our goal is to leverage NetVLAD’s permutation invariance [2] to extract rotation-invariant descriptors from rotation-equivariant feature maps, as demonstrated in [6]. This completes the final step for rotation-invariant place recognition. By concatenating the descriptors from both branches, place recognition with both accuracy and rotation-invariance can be achieved.
@@ III-C Multimodal Feature Fusion
Polar cross-attention module. In MPR, directly fusing less discriminative modalities may impair performance (see Sec. IV-E). Thus, a key challenge in LiDAR-radar fusion, especially with single-chip radar, is mitigating data sparsity and noise to uncover discriminative radar features. To extract discriminative radar features, we develop a polar cross-attention module. It captures cross-modality correlations while enhancing the discriminability of radar features through attentive fusion, as detailed in Fig. 3.
To facilitate cross-attention, we first align the dimensions of the feature maps. We first compress LiDAR feature channels through convolution. This ensures the LiDAR feature matches the radar feature in dimensions. Next, we flatten both features into sequences $\mathcal{H}_{\text{P}}\in\mathbb{R}^{B\times HW\times C_{\text{R}}}$ and $\mathcal{H}_{\text{R}}\in\mathbb{R}^{B\times HW\times C_{\text{R}}}$, which are then processed by cross-attention for multimodal fusion.
@@
[…]
where $\bar{\mathcal{A}}$ denotes the feature split of $\mathcal{A}$, $\mathrm{LN}(.)$ refers to layer normalization, $\mathcal{Q}_{\text{R}}$, $\mathcal{K}_{\text{R}}$, $\mathcal{V}_{\text{R}}$ are substantially the query, key, and value splits obtained by linearly transforming the splits of $\mathcal{H}_{\text{R}}$, and $\mathcal{Q}_{\text{P}}$, $\mathcal{K}_{\text{P}}$, $\mathcal{V}_{\text{P}}$ are those derived by the splits of $\mathcal{H}_{\text{P}}$.
Ultimately, the cross-attention enhanced radar feature is concatenated with LiDAR feature to generate the descriptor of the fusion branch. The proposed cross-attention mechanism extracts more discriminative feature representations by capturing cross-modal correlations.
Figure 3: The pipeline of the polar cross-attention module. It consists of two cross-attention branches to facilitate feature interaction of the feature map, thus exploiting cross-modality correlations.
Structure-aware distillation. The cross-modality feature fusion facilitates the extraction of discriminative multimodal place descriptions. However, while deep attentive fusion extracts discriminative features, we found it inadequate for fully exploiting radar’s robustness (see Sec. IV-E). To address this, we propose the radar-exclusive distillation branch. Furthermore, to prevent less discriminative radar descriptors from negatively impacting recognition performance, we enhance their discriminability via cross-modality distillation.
We aim to improve the performance of the distillation branch through self-distillation. The structure-aware loss [22] (formally defined in III-D) transfers geometric knowledge from the fusion branch by enforcing angular constraints on descriptor triplets. This geometric constraint ensures the descriptor triplets maintain their distributions while assimilating structural knowledge from the fusion branch. This enables the distillation branch to learn the discriminability of multimodal fusion features while exploiting the radar-specific robustness (see Sec. IV-E).